OpenStudy (anonymous):

Two end points of a line segment are (-12, -2) and (-6, -10). What are the coordinates of the point on the line through which its bisector passes?

OpenStudy (accessdenied):

What would be the actual name of the point that the bisector must pass through? Recall that the bisector divides a segment into two congruent segments, that should help.

OpenStudy (anonymous):

i tried using the midpoint formula but came out to a wrong answer (-3,4)

OpenStudy (accessdenied):

Hm, I think you may have messed up somewhere using the midpoint formula... \[ M\left( \frac{x_1 + x_2}{2} , \ \frac{y_1 + y_2}{2} \right) \] It seems like you may have subtracted instead of adding.

OpenStudy (anonymous):

after that, i got to -6/2 , 8/2. No?

OpenStudy (accessdenied):

-12 + (-6) = -12 - 6 = -18

OpenStudy (anonymous):

Ohh alright. I see the problem now.

OpenStudy (accessdenied):

Ah, okay. :) (-18/2, -12/2), or (-9, -6) is where I end up.

OpenStudy (anonymous):

Yeah i get that now haha. thank you!

OpenStudy (accessdenied):

You're welcome. :)